OpenStudy (anonymous):

PLEASE HELP!!! f(x)= sqrt(x^4-16x^2) a)Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5

sam (.sam.):

Domain in attachment

sam (.sam.):

b) Find f '(x) \[\sqrt{(x-4)}+\frac{x}{2\sqrt{x-4}}\]

OpenStudy (anonymous):

thank you very much by any chance do you know c